OpenStudy (anonymous):

A cart with mass 0.340 kg moving on a frictionless linear air track at 1.2 m/s strikes a second cart of unknown mass at rest. The collision between the two carts is elastic. After the collision, the first cart continues in its original direction at 0.66 m/s.? what is the mass of the second cart? what is the velocity of the second cart after impact?

OpenStudy (shamim):

Required formula is m1*u1+m2*u2=m1*v1+m2*v2 (1/2)*m1*u1^2+(1/2)*m2*u2^2=(1/2)*m1*v1^2+(1/2)*m2*v2^2

OpenStudy (shamim):

Response plz!!!